# Dicke superradiant enhancement of the heat current in circuit QED

### v1

@@ III Impact of a parasitic bath
The dynamics governed by the Lindblad master equation in Eq. (9) preserves the norm of the collective spin operator $\hat{J}^{2}$. Consequently, the steady-state current depends on the probabilities $P_{J}$, which are fixed by the initial state as described in Eq. (17). However, a dependence of the steady-state heat current on the initial state preparation of the qubits is not realistic in practice. Although short-term dynamics could be influenced by the initial state preparation, the long-term steady state is likely to be dominated by factors such as noise and local dissipation. These effects would eventually erase the memory of the initial state. Indeed, as we now show, such parasitic effects have a large impact on the steady-state heat current, even in the limit of vanishing small parasitic coupling, thus a non-perturbative effect. To address this effect, we introduce a third, parasitic thermal bath that interacts locally with each qubit. Unlike the collective interactions from the primary thermal baths, this parasitic bath interacts uniformly but locally with each qubit, thereby breaking the conservation of $\hat{J}^{2}$. We characterize the parasitic bath by a temperature $T_{\rm p}$ and a coupling strength $\gamma_{\rm p}$.
The Lindblad superoperator capturing the dissipative effects of the parasitic bath, $\mathcal{D}_{\rm p}(\hat{\rho})$, can be expressed as
@@
Here, $n_{\rm p}={1}/\{\exp\left[\hbar\omega_{0}/(k_{\rm B}T_{\rm p})\right]-1\}$ represents the mean thermal occupation number of the parasitic bath.
The overall dissipative dynamic of the system influenced by all three baths is thus described by
@@
Figure 5: (Color online) Panels (a,c) show the heat current as a function of the number of qubits $N$ for various values of the ratio $T_{0}/T_{\rm p}$. The independent case $\dot{Q}^{\rm ind}_{\rm h}$ is shown as a black dashed line. Panels (b,d) show the total angular momentum $\hat{J}^{2}$ (normalized by $N/2(N/2+1)$) as a function of $N$ for various values of the ratio $T_{0}/T_{\rm p}$. In panel (a,b) we have set $T_{\rm h}=T_{0}$ while $T_{\rm h}=3T_{0}$ in panel (c,d). Other parameters are $T_{\rm c}=T_{0}/3$, $\gamma_{\rm h}=\gamma_{\rm c}=\omega_{0}$, $\gamma_{\rm p}=0.001\omega_{0}$.
@@
In Fig. 5 (a,c) we show how the heat current $\dot{Q}_{\rm h}$ is influenced by the parasitic bath. Panel (a) presents a scenario where both primary thermal baths have temperatures lower than $T_{0}$, with $T_{0}/T_{\rm h}=1$ and $T_{0}/T_{\rm c}=3$. This corresponds to the system being effectively colder than $T_{0}$, $T_{0}/T^{*}\approx 1.42$. A hot parasitic bath $T_{0}/T_{\rm p}\leq 1$ suppresses the heat current even below the independent case in Eq. (25) (plotted as a black dashed line). In the particular case where $T_{\rm p}\approx T^{*}\approx 1.42T_{0}$, the heat current coincides with the value derived for independent qubits, according to Eq. (25). Under this condition, the parasitic bath has the only effect of suppressing the coherence between different qubits, rendering them effectively independent. However, when the parasitic bath is sufficiently cold $T_{0}/T_{\rm p}\geq 10$, the heat current $\dot{Q}^{\rm ind}_{\rm h}$ coincides with the one observed when the dynamics are constrained to a single subspace with $J=N/2$ ( represented by black triangles). Even in this single subspace, a super-extensive scaling in $N$ is absent. This absence is attributed to the low temperature of the baths, which maintains the system close to its ground state, leading to $\langle-\hat{J}_{z}\rangle\approx N/2$ in Eq. (24), which corresponds to a linear behavior.
@@
Panel (c) illustrates a scenario where $T_{0}/T_{\rm h}=1/3$ and $T_{0}/T_{\rm c}=3$. In this case, the system has an effective temperature bigger than $T_{0}$, $T_{0}/T^{*}\approx 0.57$. Similar to the case in panel (a), a hot parasitic bath negatively impacts the heat current. If $T_{\rm p}\approx T^{*}\approx 0.57T_{0}$, the heat current reverts to the value computed for independent atoms using Eq. (25). However, when the parasitic bath is sufficiently cold ($T_{0}/T_{\rm p}\geq 3$), the heat current exhibits super-extensive scaling with $N$. Despite this, there remains a discernible gap between the heat current $\dot{Q}^{\rm ind}_{\rm h}$ and the one calculated using Eq. (26) by setting $\bar{J}=N/2$ (represented by black triangles). This gap exists because the effective temperature of the system is relatively high $T_{0}/T^{*}\approx 0.6$, and the parasitic bath serves to redistribute the population across various subspaces characterized by different values of $J$.
In Fig. 5 (b,d) we show the dependence of the total angular momentum $\hat{J}^{2}$ on the temperature of the parasitic bath. This quantity allows us to better understand the previously discussed phenomenology of the heat current. In both panels, we see that a hot parasitic bath corresponds to low values of $\langle\hat{J}^{2}\rangle$. This can be understood as follows: a hot bath leads to a non-negligible population of multiple subspaces within the system, each characterized by different values of $J$. In the extreme case of an infinitely hot bath, all states would be equally populated. It is noteworthy that most states are concentrated around $J\approx 0$, corresponding to $\langle\hat{J}^{2}\rangle\approx 0$. In the case of high temperature, this concentration around $J\approx 0$ acts to inhibit any super-extensive scaling. Indeed, in situations where $\langle\hat{J}^{2}\rangle\approx 0$, regardless of the value of the effective temperature $T^{*}$, the heat current is reduced. This reduction can be understood using Eqs. (27,28), assuming that $\bar{J}\sim 0$ and hence does not scale with $N$. As a result, the heat current is not only diminished compared to scenarios with super-extensive scaling but is also suppressed relative to the independent case, as seen in Fig. 5 (a,c).
@@
Therefore, to achieve super-extensive scaling, a cold parasitic bath $T_{\rm p}\ll T_{0}$ is required to ensure that the subspace with maximum collective spin $J\approx N/2$ is occupied, corresponding to $\langle\hat{J}^{2}\rangle\approx N/2(N/2+1)$ in Fig. 5 (b,d). However, this requirement appears to conflict with the need for a high effective temperature $T^{*}$ for having a super-linear scaling emerging from Eq. (28). Nevertheless, given that the parasitic bath is weakly coupled to the system, it does not significantly influence the system’s effective temperature $T^{*}$, which is dominated by $T_{\text{h}}$ and $T_{\text{c}}$. As a result, a cold parasitic bath $T_{\rm p}\ll T_{0}$ does not necessarily force the system toward its ground state, allowing for a finite window where super-extensive scaling of the heat current can be observed when $T^{*}\gg T_{0}$, as shown in Fig. 5 (b).
@@ IV Experimental proposal
In this Section, we discuss an experimental proposal to observe the enhanced heat current in a circuit QED platform. In Fig. 6 (a) we plot the dependence of the heat current $\dot{Q}_{\rm h}$ on the number of qubits for different temperatures of the parasitic bath, as in Fig. 5, but this time using physical parameters - reported in Tab. 1 - that have been successfully used to describe a previous experiment Ronzani et al. (2018). This plot demonstrates that the collective enhancement of the heat current is quite pronounced when using parameters that are feasible in experimental settings.
@@
Table 1: Physical parameters compatible with the experiment in Ref. Ronzani et al., 2018. In the first line, we reported the frequencies $\nu_{T}$ associated with the relative temperatures $T$ as $\nu_{T}\equiv k_{\rm B}T/h$.
@@
Figure 6: (Color online) Panel (a) shows the heat current $\dot{Q}_{\rm h}$ (divided by $N$) and expressed in femto Watt (fW) as a function of $N$, for different values of $T_{\rm p}$ in milliKelvin (mK). Here, we choose $\omega_{0}/(2\pi)=4.0~{}{\rm GHz}$, $\gamma_{\rm h}=1.0~{}{\rm GHz},\gamma_{c}=1.0~{}{\rm GHz}$. Panel (b) shows the heat current $\dot{Q}_{\rm h}$ in femto Watt (fW) as a function of $T_{\rm p}$ in milliKelvin (mK) for $N=2$ qubits. In this panel, the collective heat current $\dot{Q}_{\rm h}$ is depicted by a blue continuous line, the heat current in the independent scenario $\dot{Q}^{\rm ind}{\rm h}$ is represented by a red dashed line, and the heat current in the detuned scenario is indicated by a green dotted line. For the detuned scenario, the parameters are set as $\omega^{(1)}_{0}/(2\pi)=4.0~{}{\rm GHz}$ and $\omega^{(2)}_{0}/(2\pi)=2.0~{}{\rm GHz}$, with $\gamma_{1}=1.0~{}{\rm GHz}$. The rate $\gamma_{2}$ is calculated in accordance with Eq. (37), assuming $\mathcal{Q}_{i}=20$. In both panels, additional parameters are $\nu_{T_{\rm h}}=8.0~{}{\rm GHz}$, $\nu_{T_{\rm c}}=1.0~{}{\rm GHz}$ (corresponding to bath temperatures $T_{\rm h}=384.0~{}{\rm mK}$ and $T_{\rm c}=96.0~{}{\rm mK}$), and a parasitic bath coupling rate of $\gamma_{\rm p}=0.01~{}{\rm GHz}$.
@@
Furthermore, we propose a minimal single-device experiment to demonstrate the collective advantage in a setup with $N=2$ transmon qubits. Ideally, we would like to compare the collective current $\dot{Q}_{\rm h}$ with the independent case $\dot{Q}^{\rm ind}_{\rm h}$ in Eq. (25), where each qubit is independently coupled with its bath. However, working with two different devices introduces variability in Hamiltonian parameters due to fabrication differences which makes a direct comparison difficult. To circumvent this problem, we devise a protocol where we exploit the possibility of controlling the frequency of each transmon by adjusting the external magnetic flux threading through it. By controlling the detuning between the qubits we can move in practice from the collective to the independent coupling scheme. Let us therefore assume that the two transmon qubits have tunable frequencies, $\omega_{0}^{(1)}$ and $\omega_{0}^{(2)}$. When both of these frequencies are resonant with the LC resonator frequency ($\omega_{0}^{(k)}=\omega_{\rm LC}$ for $k=1,2$), the transmons are collectively coupled to the same modes of the baths. Under this condition, the heat current is the collective one $\dot{Q}_{\rm h}$, determined by Eq. (22) and discussed in previous sections. However, if we detune e.g. the second qubit setting $\omega_{0}^{(2)}=\omega_{\rm LC}/2$, while keeping the first qubit on resonance $\omega_{0}^{(1)}=\omega_{\rm LC}$, we can effectively “disconnect” the second qubit.
We label the heat current in this configuration as $\dot{Q}^{\rm det}_{\rm h}$. In this detuned scenario, the qubits interact with independent modes of the baths and behave independently. Additionally, due to the significant detuning of the second qubit, it contributes negligibly to the heat current. Therefore, we expect that only the first qubit contributes significantly to the heat current, hence we expect that $\dot{Q}^{\rm det}_{\rm h}\approx\dot{Q}^{\rm ind}_{\rm h}/2$.
We now detail the calculation of the heat current $\dot{Q}^{\rm det}_{\rm h}$ in this detuned scenario, and quantitatively validate the analysis above. In this case, each qubit interacts independently with distinct modes of the bath. The total heat current is thus given by the sum of the independent contribution of each qubit, each one determined by Eq. (22) setting $N=1$, with the qubit’s state determined by the master equation in Eq. (32). We assume equal strength in the couplings with both the hot and cold baths. Therefore, in the detuned scenario, the total heat current in Eq. (22) can be expressed as:
$\begin{split}\dot{Q}^{\rm det}_{\rm h}&=\sum_{k=1,2}\hbar\omega_{0}^{(k)}\gamma_{k}\Big{[}-(1+n^{(k)}_{\rm h})\langle\hat{\sigma}^{(k)}_{+}\hat{\sigma}^{(k)}_{-}\rangle+\\ &+n^{(k)}_{\rm h}\langle\hat{\sigma}^{(k)}_{-}\hat{\sigma}^{(k)}_{+}\rangle\Big{]}~{},\end{split}$
where $\gamma_{k}$ represents the decay rate of the $k$-th qubit and $\hat{\sigma}^{(k)}_{-}$ ($\hat{\sigma}^{(k)}_{+}$) are the creation (destruction) Pauli operators acting on the $k$-th qubit and the thermal occupations $n^{(k)}_{i}$ of the $i$-th bath depend on the bath temperature $T_{i}$ and the frequency of the $k$-th qubit:
$\langle\hat{\sigma}^{(k)}_{+}\hat{\sigma}^{(k)}_{-}\rangle$ $=p^{(k)}_{1}~{},$ (34)
$\langle\hat{\sigma}^{(k)}_{-}\hat{\sigma}^{(k)}_{+}\rangle$ $=p^{(k)}_{0}~{},$
Hence, Eq. (IV) can be recasted as:
$\gamma_{2}=\frac{\gamma_{1}}{{1+\mathcal{Q}_{i}^{2}({\omega_{0}}/{\omega_{\rm LC}}-{\omega_{\rm LC}}/{\omega_{0}})^{2}}}~{},$ (37)
with $\mathcal{Q}_{i}$ being the quality factors of the resonators.
In Fig. 6 (b), we show the heat current as a function of the temperature of the parasitic bath for the minimal single-device system comprising $N=2$ transmon qubits. The continuous blue line represents the collective heat current $\dot{Q}_{\rm h}$, while the red dashed line indicates the heat current in the independent case $\dot{Q}^{\rm ind}_{\rm h}$, calculated as per Eq. (25). The green dotted line depicts twice the heat current in the scenario where one qubit is in resonance and the other is detuned, namely $2\dot{Q}^{\rm det}_{\rm h}$ and determined by Eq. (36). As we can see in the figure, the detuned case can be used to estimate the independent scenario, as $2\dot{Q}^{\rm det}_{\rm h}\approx\dot{Q}^{\rm ind}_{\rm h}$.
Significantly, Fig. 6 (b) reveals that at a parasitic bath temperature of $T_{\rm p}=50$ mK, the collective $\dot{Q}_{\rm h}$ exhibits an enhancement of approximately $13\%$ compared to the independent scenario. In contrast, at $T_{\rm p}=450$ mK, the collective heat current shows a reduction of about $4\%$ relative to the independent case.

### v3

@@ III Impact of a parasitic bath
The dynamics governed by the Lindblad master equation in Eq. (9) preserves the norm of the collective spin operator, $\hat{J}^{2}$. Consequently, the steady-state current depends on the probabilities $P_{J}$, which are fixed by the initial state as described in Eq. (17). However, a dependence of the steady-state heat current on the initial state preparation of the qubits is not realistic in practice. Although short-term dynamics could be influenced by the initial state preparation, the long-term steady state is likely to be dominated by factors such as noise and local dissipation. These effects would eventually erase the memory of the initial state. Indeed, as we now show, such parasitic effects have a large impact on the steady-state heat current, even in the limit of vanishing small parasitic coupling, thus a non-perturbative effect. To address this effect, we introduce a third, parasitic thermal bath that interacts locally with each qubit. Unlike the collective interactions from the primary thermal baths, this parasitic bath interacts uniformly but locally with each qubit, thereby breaking the conservation of $\hat{J}^{2}$. This parasitic bath is characterized by a temperature $T_{\rm p}$ and a coupling strength $\gamma_{\rm p}$.
The dissipator associated with the parasitic bath, $\mathcal{D}_{\rm p}(\hat{\rho})$, can be expressed as
@@
Here, $n_{\rm p}={1}/\{\exp\left[\hbar\omega_{0}/(k_{\rm B}T_{\rm p})\right]-1\}$ represents the mean thermal occupation number of the parasitic bath. The overall dissipative dynamic of the system influenced by all three baths is thus described by
@@
Figure 4: (Color online) Panels (a,c) show the heat current, divided by $N$, as a function of the number of qubits $N$ for various values of the ratio $T_{0}/T_{\rm p}$. The independent case $\dot{Q}^{\rm ind}_{\rm h}$, given by Eq. (26), is shown as a black dashed line. The ideal case without the parasitic bath, and calculated assuming Eq. (27) with $\bar{J}=N/2$, is shown as black triangles. Panels (b,d) show the total angular momentum $\hat{J}^{2}$ (normalized by $N/2(N/2+1)$) as a function of $N$ for various values of the ratio $T_{0}/T_{\rm p}$. In panel (a,b) we have set $T_{\rm h}=T_{0}$ while $T_{\rm h}=3T_{0}$ in panel (c,d). Other parameters are $T_{\rm c}=T_{0}/3$, $\gamma_{\rm h}=\gamma_{\rm c}=\omega_{0}/10$, $\gamma_{\rm p}=0.001\omega_{0}$.
@@
In Fig. 4 (a,c) we show how the heat current $\dot{Q}_{\rm h}$, divided by $N$, is influenced by the parasitic bath. Panel (a) focuses on the low-temperature regime, i.e. when both primary thermal baths have temperatures lower than $T_{0}$, with $T_{0}/T_{\rm h}=1$ and $T_{0}/T_{\rm c}=3$. This corresponds to the system being effectively colder than $T_{0}$, $T_{0}/T^{*}\approx 1.42$. As we can see, a hot parasitic bath $T_{0}/T_{\rm p}\leq 1$ (blue and orange dots) suppresses the heat current even below the independent case in Eq. (26) (plotted as a black dashed line). In the particular case where $T_{\rm p}\approx T^{*}\approx 1.42T_{0}$ (green dots), the heat current coincides with the value derived for independent qubits, according to Eq. (26). Under this condition, the parasitic bath has the only effect of suppressing the coherence between different qubits, rendering them effectively independent. However, when the parasitic bath is sufficiently cold $T_{0}/T_{\rm p}\geq 10$ (red and purple dots), the heat current $\dot{Q}^{\rm ind}_{\rm h}$ coincides with the one observed when the dynamics are constrained to a single subspace with $J=N/2$ ( represented by black triangles). Even in this single subspace, a super-extensive scaling in $N$ is absent. This absence is attributed to the low-temperature regime of the baths, which maintains the system close to its ground state, leading to $\langle-\hat{J}_{z}\rangle\approx N/2$ in Eq. (24), which corresponds to a linear behavior.
@@
Panel (c) illustrates the high-temperature scenario where $T_{0}/T_{\rm h}=1/3$ and $T_{0}/T_{\rm c}=3$. In this case, the system has an effective temperature bigger than $T_{0}$, $T_{0}/T^{*}\approx 0.57$. Similar to the case in panel (a), a hot parasitic bath (blue dots) negatively impacts the heat current. If $T_{\rm p}\approx T^{*}\approx 0.57T_{0}$ (orange dots), the heat current reverts to the value computed for independent atoms using Eq. (26). However, when the parasitic bath is sufficiently cold ($T_{0}/T_{\rm p}\geq 3$, red and purple dots), the heat current exhibits super-extensive scaling with $N$.
Despite this, a discernible gap remains between the heat current $\dot{Q}_{\rm h}$ calculated in the presence of the parasitic bath, and the ideal one without the parasitic bath calculated assuming Eq. (27) with $\bar{J}=N/2$ (represented by black triangles). This gap exists because the parasitic bath redistributes the population across various subspaces characterized by different values of $J$, whereas in Eq. (27) we assume that a single value of $\bar{J}=N/2$ is occupied.
To further interpret our results, in Fig. 4 (b,d) we show the dependence of the total angular momentum $\langle\hat{J}^{2}\rangle$, divided by $\sim N^{2}$, on the temperature of the parasitic bath. In both panels, we see that a hot parasitic bath (blue, orange, and green dots) corresponds to low values of $\langle\hat{J}^{2}\rangle$, which is consistent with low values of the heat currents shown in the corresponding panels (a) and (c). This can be interpreted in the following way. Let us consider the high-temperature regime of the main baths, which is the one that can exhibit the super-linear scaling of the heat current (see Sec. II.3). In the limit of infinitely hot parasitic bath, all states would be equally populated. As $N$ increases, the states with $J\sim 0$ exponentially outnumber the other states, leading to the observed decrease of $\hat{J}^{2}$. This, in turn, implies a vanishing $\langle\hat{J}_{z}\rangle$, which, using Eq. (24), implies a vanishing heat currents even below the independent qubits case - as observed e.g. in the orange and blue dots in panel (a) and (c).
@@
Therefore, to achieve super-extensive scaling, a cold parasitic bath $T_{\rm p}\ll T_{0}$ is required to ensure that the subspace with maximum collective spin $J\approx N/2$ is occupied, corresponding to $\langle\hat{J}^{2}\rangle\approx N/2(N/2+1)$ in Fig. 4 (b,d). However, this requirement appears to conflict with the need for a high effective temperature $T^{*}$ to have the super-linear scaling emerging from Eq. (30). Nevertheless, given that the parasitic bath is weakly coupled to the system, it does not significantly influence the system’s effective temperature $T^{*}$, which is dominated by $T_{\text{h}}$ and $T_{\text{c}}$. As a result, a cold parasitic bath $T_{\rm p}\ll T_{0}$ does not necessarily force the system toward its ground state, allowing for a finite window where super-extensive scaling of the heat current can be observed when $T^{*}\gg T_{0}$, as shown in Fig. 4 (b).
@@ IV Experimental proposal
[…]
@@
Table 1: Physical parameters compatible with the experiment in Ref. 28. In the first line, we reported the frequencies $\nu_{T}$ associated with the relative temperatures $T$ as $\nu_{T}\equiv k_{\rm B}T/h$.
@@
Figure 5: (Color online) Panel (a) shows the heat current $\dot{Q}_{\rm h}$ (divided by $N$) and expressed in femto Watt (fW) as a function of $N$, for different values of $T_{\rm p}$ in milliKelvin (mK). Here, we choose $\omega_{0}/(2\pi)=4.0~{}{\rm GHz}$, $\gamma_{\rm h}=\gamma_{c}=0.5~{}{\rm GHz}$. Panel (b) shows the heat current $\dot{Q}_{\rm h}$ in femto Watt (fW) as a function of $T_{\rm p}$ in milliKelvin (mK) for $N=2$ qubits. In this panel, the collective heat current $\dot{Q}_{\rm h}$ is depicted by a blue continuous line, the heat current in the independent scenario $\dot{Q}^{\rm ind}_{\rm h}$ is represented by a red dashed line, and the heat current in the detuned scenario is indicated by a green dotted line. For the detuned scenario, the parameters are set as $\omega^{(1)}_{0}/(2\pi)=4.0~{}{\rm GHz}$ and $\omega^{(2)}_{0}/(2\pi)=2.0~{}{\rm GHz}$, with $\gamma_{1}=0.5~{}{\rm GHz}$. The rate $\gamma_{2}$ is calculated in accordance with Eq. (92), assuming $\mathcal{Q}_{i}=20$. In both panels, additional parameters are $\nu_{T_{\rm h}}=8.0~{}{\rm GHz}$, $\nu_{T_{\rm c}}=1.0~{}{\rm GHz}$ (corresponding to bath temperatures $T_{\rm h}=384.0~{}{\rm mK}$ and $T_{\rm c}=96.0~{}{\rm mK}$), and a parasitic bath coupling rate of $\gamma_{\rm p}=0.01~{}{\rm GHz}$.
@@
Furthermore, we propose a minimal single-device experiment to demonstrate the collective advantage in a setup with $N=2$ transmon qubits. Ideally, we would like to compare the collective current $\dot{Q}_{\rm h}$ with the independent case $\dot{Q}^{\rm ind}_{\rm h}$ in Eq. (26), where each qubit is independently coupled with its bath. However, working with two different devices introduces variability in Hamiltonian parameters due to fabrication differences which makes a direct comparison difficult. To circumvent this problem, we devise a protocol where we exploit the possibility of controlling the frequency of each transmon by adjusting the external magnetic flux threading through it. By controlling the detuning between the qubits, we can switch between the collective and the independent coupling scheme on the same device.
Let us assume that the two transmon qubits have tunable frequencies, $\omega_{0}^{(1)}$ and $\omega_{0}^{(2)}$. When both of these frequencies are resonant with the $LC$ resonator frequency ($\omega_{0}^{(k)}=\omega_{LC}$ for $k=1,2$), the transmons are collectively coupled to the same modes of the baths. Under this condition, the heat current is the collective one $\dot{Q}_{\rm h}$ determined by Eq. (22) and discussed in previous sections. However, if we detune e.g. the second qubit setting $\omega_{0}^{(2)}=\omega_{\rm LC}/2$, while keeping the first qubit on resonance $\omega_{0}^{(1)}=\omega_{LC}$, we can effectively “disconnect” the second qubit. We label the heat current in this configuration as $\dot{Q}^{\rm det}_{\rm h}$. In this detuned scenario, the qubits interact with independent modes of the baths and behave independently. Additionally, due to the significant detuning of the second qubit, it contributes negligibly to the heat current. Therefore, only the first qubit contributes significantly to the heat current, hence we expect that $\dot{Q}^{\rm det}_{\rm h}\approx\dot{Q}^{\rm ind}_{\rm h}/2$. In Appendix D, we detail the calculation of the heat current $\dot{Q}^{\rm det}_{\rm h}$ in the detuned scenario. This derivation allows us to quantitatively support the previous analysis. In Fig. 5 (b), we show the heat current as a function of the temperature of the parasitic bath for the minimal single-device system comprising $N=2$ transmon qubits. The continuous blue line represents the collective heat current $\dot{Q}_{\rm h}$, while the red dashed line indicates the heat current in the independent case $\dot{Q}^{\rm ind}_{\rm h}$, calculated as per Eq. (26). The green dotted line depicts twice the heat current in the scenario where one qubit is in resonance and the other is detuned, namely $2\dot{Q}^{\rm det}_{\rm h}$ and determined by Eq. (91). As we can see in the figure, the detuned case can be used to accurately estimate the independent scenario as $2\dot{Q}^{\rm det}_{\rm h}\approx\dot{Q}^{\rm ind}_{\rm h}$, and thus to experimentally validate and quantify the heat current enhancement in the collective case.
Significantly, Fig. 5 (b) reveals that at a parasitic bath temperature of $T_{\rm p}=50$ mK, the collective $\dot{Q}_{\rm h}$ exhibits an enhancement of approximately $13\%$ compared to the independent scenario. In contrast, at $T_{\rm p}=450$ mK, the collective heat current shows a reduction of about $4\%$ relative to the independent case. Concerning the experimental implementation, fabricating a multi-qubit circuit is not significantly more demanding than making a single qubit. We believe that the main experimental challenge in realizing such a minimal single-device experiment is ensuring that the two transmons have the same frequency $\omega_{0}$ and are coupled to the same bath with the same intensity to guarantee the conservation of $\hat{J}^{2}$. Small imperfections in the physical implementation would inevitably result in breaking the conservation of $\hat{J}^{2}$. However, heuristically, these effects would be similar to the presence of the parasitic baths that we studied earlier.